OpenStudy (anonymous):

would this x^3x^-5/x^-4 become this :x^-4/2x^2 ? if it was simplified

jimthompson5910 (jim_thompson5910):

x^3 * x^-5 = x^(3 + (-5) ) = x^(-2) So x^3 * x^-5 = x^(-2)

jimthompson5910 (jim_thompson5910):

which means x^3x^-5/x^-4 becomes x^-2/x^-4

jimthompson5910 (jim_thompson5910):

x^-2/x^-4 = x^(-2 - (-4) ) = x^(-2 + 4) = x^2 So the final answer is x^2

OpenStudy (anonymous):

wow you made it look easy :) thanks

jimthompson5910 (jim_thompson5910):

you're welcome, I'm glad it's making sense

jimthompson5910 (jim_thompson5910):

The general rule is this

jimthompson5910 (jim_thompson5910):

When multiplying exponential expressions with the same base, you add the corresponding exponents.

jimthompson5910 (jim_thompson5910):

When dividing exponential expressions with the same base, you subtract the corresponding exponents.

OpenStudy (anonymous):

oh i see where i made those mistakes thanks for the rules

jimthompson5910 (jim_thompson5910):

sure thing